OpenStudy (heart_offire):

i need help 4x + 3y = 6 -4x + 2y = 14 Solve the system of equations. A) x = 1/2, y = 3 B) x = 3, y = 1/2 C) x = 4, y = -3/2 D) x = -3/2, y = 4

OpenStudy (anonymous):

Do you know how to solve system of equations at all?

OpenStudy (heart_offire):

umm sometimes, i would plug them in but i dont like plugging them in as fraction

OpenStudy (anonymous):

Have you heard of the elimination method for solving system of equations? xD You can do it with this one. Basically, you eliminate one variable to solve for another by adding/subtracting. So, let's look at the x and y values. 4x and -4x 3y and 2y Which of these cancels each other out?

OpenStudy (heart_offire):

4x and -4x

OpenStudy (anonymous):

Right. So you can eliminate those and solve for y. I'll set the problem up and leave you to it. :D 4x + 3y = 6 -4x + 2y = 14 -------------- 0 + 5y = 20 5y = 20 (now you can solve for y) Once you obtain y, plug it into one of the equations and then you'll be able to find x.

OpenStudy (heart_offire):

y=4

OpenStudy (anonymous):

Mhm! Plug that into one of the equations. You can already see that the answer is D because it's the only answer choice where y = 4, but it might be beneficial if you finish solving.

OpenStudy (heart_offire):

when i plug it in for 4 in for y will x be 1

rebeccaxhawaii (rebeccaxhawaii):